OpenStudy (hihi67):

With the help of a bar, the height of trees or buildings can be approximately determined. It stands behind the bar on so that its upper end coincides with the peak of the building. A second or a second measures the horizontal distance between the eyes bar (a), the distance of the bar to the building (b), the distance between the eyes from the ground (d) and the total height (h) of the crossbar. Place with the variables a, b, c, x is a ratio equation! Start here with x! HELP!!!

OpenStudy (hihi67):

i can't figure out the thing with x, i mean there isn't really anything that is "like it"

OpenStudy (anonymous):

it's about the same angles here, h:c (c=building height minus the d) as a:(a+b) building height = d + h/((a+b)/a)

OpenStudy (hihi67):

umm c is not the building height. c is h -d

OpenStudy (hihi67):

x is the building height

OpenStudy (hihi67):

sort of

OpenStudy (anonymous):

ok, let me rephrase that than: h:(x-d) as a: (a+b) x = d + h/(a+b)/a I presume that the measurement stick does not reach to the ground, but is measured from eye level up.

OpenStudy (hihi67):

yes

OpenStudy (hihi67):

still confused

OpenStudy (hihi67):

and when you say x-d, do you mean....... (wait)

OpenStudy (anonymous):

The triangle you make between your eyes and the stick is similar in shape as the large triangle you can draw between your eye and the building. That's the reason that you have similar ratios for height and length of the triangles.

OpenStudy (hihi67):



OpenStudy (hihi67):

or

OpenStudy (hihi67):